Draw triangle xyz with vertices x(-3,2),Y(4,-2), and z(-3,-2) on the coordinate plane.Then find the area of the triangle in square units
1
Expert's answer
2019-12-18T12:03:17-0500

"Area=\frac{1}{2}|x_1(y_2-y_1)+x_2(y_3-y_1)+x_3(y_1-y_2)|"

Plugging values,

"x_1=-3,x_2=4,x_3=-3"

"y_1=2,y_2=-2,y_3=-2"

Plugging the values in above formula,

"Area=\frac{1}{2}|(-3)[-2-(-2)]+4(-2-2)+(-3)[2-(-2)]|"

"Area=\frac{1}{2}|(-3)[-2+2]+4(-2-2)+(-3)[2+2]|"

"Area=\frac{1}{2}|(-3)[0]+4(-4)+(-3)[4]|"

"Area=\frac{1}{2}|0-16-12|"

"Area=\frac{1}{2}|-28|"

"Area=\frac{1}{2} \times{28}"

"Area=14"

So, area = 14 square units.